How many milliliters of 3 F KOH can be prepared from 100 grams of KOH?


1
Expert's answer
2021-03-04T06:08:18-0500

According to formality definition:

F = n / V = m / (Mr × V)

where F - formality, m - mass, V - volume, molecular weight

From here:

V = n / F = m / (F × MrKOH) = 100 g / (3 F × 56 g/mol) = 0.6 L = 600 mL


Answer: 600 mL